Two angles are supplementary. Three times one angle is equal to the other angle plus the product of 8 and 9. What are the 2 angles?
Answer by jgr45(31)


First we need to translate what we're told into equation form. I'll call the angles x and 180-x (supplementary angles total 180):
3x = (180 - x) + 72 (product of 8 and 9)
First remove the grouping symbols:
3x = 180 - x + 72
Combine the constants:
3x = 252 - x
Add x to eliminate it from the right side:
4x = 252
Divide by 4:
x = 63
This is one angle; the other is its difference from 180, or 117.
So your supplementary angles are 117 and 63.
Check:
3(63) = (180 - 63) + 72
189 =? 117+72
189 = 189 check